The utility function is u = x1½ + x2, and the budget constraint is m = p1x1 + p2x2.

  1. Derive the optimal demand curve for good 1, x1(p1, p2), and good 2, x2(m, p1, p2).

  1. Looking at the cross price effects (∂x1/∂p2 and ∂x2/∂p1) are goods x1 and x2 substitutes or complements? Looking at income effects (∂x1/∂m and ∂x2/∂m) are goods x1 and x2 inferior, normal or neither?

  1. Assume m=100, p1=0.5 and p2=1. Using the demand function you derived in part a, what is the consumption of x1 and of x2?

  1. Consider a price drop for only good 1 to p1’=0.25. Calculate the new demands for good 1 and good 2 and label them x1M and x2M. Plot in a graph with p1 on the vertical axis the x1 you found in part c, and x1M. Graph the resulting Marshallian demand curve.

  1. Assume you are endowed with the amounts x1 and x2 you found in part c. What is their total worth, given the new prices? Define this as mS = p1’x1 + p2x2. If mS was your original budget constraint and you had to maximize your utility, what would be the resulting optimal consumption x1S(p1’,p2) and x2S(mS,p1’,p2)?

   f. Plot the original x1 from part c and x1S from part e in a graph with p1 on the vertical axis. Draw the Slutsky demand curve. How does it differ from the Marshallian demand curve? What is the substitution effect and the income effect of this price change?
Answer #1

It's mandatory to answer only first 4 parts

& UCx1, x2) = NX1 + X2 = Pixi + P2X2 a) at eam MRS112 = PilP2 MRS=MUL = 1 x at Eqm ax = 1/2 = x1 = (P2) 2 MO 2 from B.C. x2 *(15) ad curve (4,•25) Xim.